OpenStudy (iwatermelon):

Which ordered pairs are solutions to the inequality 4x+y>−64x+y>−6? Select each correct answer. (4, −20)(4, −20) (−3, 6)(−3, 6) (−1, −1)(−1, −1) (2, 0)(2, 0) (0, −9)

OpenStudy (jango_in_dtown):

4x+y>−64x+y implies -68x<0 implies x>0

OpenStudy (iwatermelon):

?

OpenStudy (jango_in_dtown):

can you provide a snapshot of the question?

OpenStudy (iwatermelon):

all of it is up there

OpenStudy (iwatermelon):

i dont know how to snap shot xd

OpenStudy (jango_in_dtown):

ok so we got x>0

OpenStudy (jango_in_dtown):

solutions can be from the first and 4th options only

OpenStudy (iwatermelon):

oh ok thanks

OpenStudy (jango_in_dtown):

also see 4x+y>-6 put x=2 and y=0 , see it is satisfied so (2,0) is solution

OpenStudy (iwatermelon):

ok

OpenStudy (jango_in_dtown):

put x=4, y=-20 then 4x+y=-16-20=-36<-6, so (4,-20) is not a solution

OpenStudy (jango_in_dtown):

hence only one solution (2,0)

OpenStudy (iwatermelon):

ok thanks gtg